Question 5

Michael and Janine are both investors wanting to purchase investment portfolios. The expected return, standard deviation and beta of various investments are shown in the following table:

Investment Average
return over last year (% pa)
Standard
Deviation (% pa)
Beta
Apocryphal Industries 19 13 2.0
Bona Fide Ltd 10 8 0.7
Market Index 12 2 1
10 year Bond 5 2 0

Michael wants a portfolio that has an expected return of 12% pa and that contains only Apocryphal Industries and 10 year bonds.

Janine wants a portfolio with a 75% weighting in Apocryphal Industries and a 25% weighting in Bona Fide Ltd shares. The correlation coefficient between the returns of the shares in Apocryphal Industries and Bona Fide Ltd is 0.6.

a)Calculate the percentage of each share that Michael should include in his portfolio to achieve his desired results. Give your answers as a percentage to the nearest percent.

Investment Percentage to invest
Apocryphal Industries % pa
Bona Fide Ltd % pa
10 year Bond % pa

b)Janine is interested in finding out how risky her portfolio is. Calculate the standard deviation of Janine's portfolio. Give your answer as a percentage to 2 decimal places.

Standard deviation =  % pa

Question 6

You own a portfolio that comprises 2 assets:

  • 1,000 shares of company X each valued at $5 with an expected return of 8% pa
  • 1,500 shares of company Y each valued at $2 with an expected return of 12% pa

Calculate the expected return of the portfolio. Give your answer as a percentage per annum to 1 decimal place.

Expected return =  % pa
Answer #1

1.
w*19%+(1-w)*5%=12%
=>w=0.5

Apocryphal 50%
Bona Fide 0%
10 year Bond 50%

2.
=sqrt((75%*13%)^2+(25%*8%)^2+2*75%*25%*13%*8%*0.6)=11.0663%

3.
=(1000*5*8%+1500*2*12%)/(1000*5+1500*2)=9.5000%